linear algebra

(3) Find the QR factorization of A.

T 2 31 2 4 1,b= 1 1 1 [1] 2 1 Given: A =

Thank you for the help.
Answer #1

[ 1 3 A = - NN 4 4 Let Ag = [ 2, 2, 1], Q2 = [ 3, 4, 1]. Now, perform the Gram-Schmidt procedure Ug = Q4 -. [ 2, 2, 1] est10 4 Uz = [3, 4, 1] - 16+8+1).} } [2, 2, 1] - [ 3, 4, 1]- $ [ 2; 2, 1] = (3-40, 4-10, 1-3] ; U = [- £, £, - }] = { [-3, 2Qg. @s = { [ 22+2° +12) = 3 az: es = {( 3x2 + 4x2 +1x 1) = f(6+8 + 1) = 5 az.bz = { [3X(-1) + 4 x2 +1 X(-2)] = { 1-3 +8-2) =2 2 1 50. Rx = 1 1 1 2X1 + 2X2 + 1 X 1 حالي -1X1 + 2X2 - 2X1 ا 77 2+4+1 حولي) 1 -1 +4-2 RX = WA or 3 5 7 127 10 1] [22] حالي)